OpenStudy (anonymous):

Mercury has a mass of 3.30 1023 kg and a radius of 2.44 106 m. (a) What is the acceleration due to gravity on Mercury?

OpenStudy (anonymous):

acceleration = G * M/R^2 G = 6.67 * 10^-11 N(m/kg)^2 M - mass of the planet R - radius